# Predictive Modeling: BIM Command Recommendation Based on Large-Scale Usage Logs

### v1

@@ Abstract
The adoption of Building Information Modeling (BIM) and model-based design within the Architecture, Engineering, and Construction (AEC) industry has been hindered by the perception that using BIM authoring tools demands more effort than conventional 2D drafting. To enhance design efficiency, this paper proposes a BIM command recommendation framework that predicts the optimal next actions in real-time based on users’ historical interactions. We propose a comprehensive filtering and enhancement method for large-scale raw BIM log data and introduce a novel command recommendation model. Our model builds upon the state-of-the-art Transformer backbones originally developed for large language models (LLMs), incorporating a custom feature fusion module, dedicated loss function, and targeted learning strategy. In a case study, the proposed method is applied to over 32 billion rows of real-world log data collected globally from the BIM authoring software Vectorworks. Experimental results demonstrate that our method can learn universal and generalizable modeling patterns from anonymous user interaction sequences across different countries, disciplines, and projects. When generating recommendations for the next command, our approach achieves a Recall@10 of approximately 84%.
@@
[inst1] organization=Chair of Computational Modeling and Simulation, School of Engineering and Design, Technical University of Munich,country=Germany
@@ 2.2 Transformers for sequential recommendation
Recent advancements in large language models (LLMs) have highlighted their remarkable ability to comprehend language sequences. The underlying Transformer architecture, originally developed for Natural Language Processing (NLP) tasks, has gained widespread adoption across various fields due to its scalability and proficiency in modeling complex patterns in sequential data. The sequential nature of user interactions aligns closely with language modeling tasks, leading to the adoption of NLP-inspired architectures in recommendation systems [22, 23, 24]. The success of the Transformer architecture lies in its attention mechanism, which effectively captures dependencies between representations regardless of their position in the sequence [22]. This capability makes it particularly well-suited for modeling the dynamic and evolving behavior of users – akin to the dynamic design process where designers’ actions and interests vary across workflows and projects [1].
Transformer-based recommendation systems often enhance user interaction sequences by incorporating meta-information as additional features. In the e-commerce scenario, such features may include user comments, product descriptions, prices, or images [25]. This differs from the Transformers in NLP, where additional meta-information is absent. Language models typically consist of three components: (1) a tokenizer that converts raw text into a sequence of index-encoded tokens, (2) a Transformer architecture that learns latent representations of the sequence, and (3) a task-specific prediction head designed for applications such as sentiment analysis or next-token prediction [24].
In our study, we focus exclusively on the (2) Transformer architecture to model the sequential dependencies among commands. We replace the tokenizer and NLP-specific prediction heads with a feature fusion module capable of integrating additional information and a prediction head tailored for the command recommendation task. Our implementation builds on Transformer4Rec [24], a flexible open-source framework that bridges NLP and sequential recommendation systems. By leveraging the Transformer backbone from language models like XLNet [26], this framework has demonstrated state-of-the-art performance in generating recommendations for news and e-commerce domains. However, our research specifically explores Transformers for BIM command recommendation, which poses unique challenges such as the lack of rich meta-information, varying session lengths, and the long-tail distribution of commands.
@@ 3.1 Data filtering and enhancement
Previous studies have primarily focused on developing data processing pipelines based on Revit logs; a comprehensive review is provided in [5]. However, these studies have not addressed more challenging issues related to logs from other BIM authoring software, which often exhibits multilingual content, a mix of high- and low-level commands, and excessive complexity. As an example, Figure 3 illustrates a snippet of the Vectorworks log containing two sessions in different languages.
@@ 3.1.1 Actual modeling flow tracking
Algorithm 1 Matching and Removal of Redo and Undo Commands
1:Initialize $recent\_commands\leftarrow\emptyset$
2:Initialize $recent\_undone\_commands\leftarrow\emptyset$
3:Initialize $to\_remove\_indices\leftarrow\emptyset$
4:for each $(\text{cmd},idx)$ in each log session do
5:    if cmd is a normal command then
6:        Append $(\text{cmd},idx)$ to $recent\_commands$
[…]
8:        Search backward in $recent\_commands$ for a matching command $(c,i)$
9:        if found then
10:           Add $idx$ and $i$ to $to\_remove\_indices$
11:           Remove $(c,i)$ from $recent\_commands$
12:           Append $(c,i)$ to $recent\_undone\_commands$
13:        end if
[…]
16:        if found then
17:           Remove $i$ from $to\_remove\_indices$
18:           Add $idx$ to $to\_remove\_indices$
19:           Remove $(c,i)$ from $recent\_undone\_commands$
20:           Append $(c,i)$ to $recent\_commands$
21:        else
22:           Add $idx$ to $to\_remove\_indices$
23:        end if
24:    end if
25:end for
@@ 3.1.2 Multi-language alignment
If the similarity falls below the threshold, we assume that the ID is shared by multiple different commands, as the English names obtained by the translation API show significant semantic differences. In such cases, the unsupervised clustering algorithm DBSCAN [28][29] is employed to further group the commands into sub-clusters. Each sub-cluster is subsequently labeled using the command name of its midpoint vector. The process is outlined in Alg. 2. The example outcomes from this module are illustrated in Table 1.
Algorithm 2 Multi-language Alignment
1:Actual modeling logs S
2:Predefined similarity threshold $\phi$
3:Aligned logs $\textbf{S}_{\textit{aligned}}$
6:for $\textbf{S}_{\textit{id}}\in\textbf{S}_{\textit{id\_groups}}$ do
7:    for $\text{s}\in\textbf{S}_{\textit{id}}$ do
10:    end for
12:    if $\phi_{\textit{id}}\geq\phi$ then
15:    else
20:        end for
21:    end if
22:end for
25:return $\textbf{S}_{\textit{aligned}}$
@@ 3.1.3 Redundant command identification and removal
Algorithm 3 Redundant Command Identification
1:Aligned logs $\textbf{S}_{\textit{aligned}}$
2:Similarity threshold $\phi$
3:High-/low-level command mapping $\textbf{S}_{\textit{mapping}}$
4:$\textbf{S}_{\textit{mapping}}\leftarrow\emptyset$
7:    $\textbf{S}_{\textit{pairs}}\leftarrow\emptyset$
[…]
13:        if $\text{confidence}>\phi$ then
15:           if approved then
17:           end if
18:        end if
19:    end for
20:    if $\textbf{S}_{\textit{pairs}}\neq\emptyset$ then
22:    end if
23:end for
24:return $\textbf{S}_{\textit{mapping}}$
@@ 3.2.5 Multi-task learning and loss function
where $p_{t}$ is the model’s estimated probability for the ground-truth class, $\alpha_{t}$ is a weighting factor to address class imbalance, and $\gamma$ is a focusing parameter that smoothly adjusts the rate at which easy examples are down-weighted. Let the losses for the three tasks be $\ell_{\text{cmd}},\ell_{\text{cls}},\ell_{\text{tgt}}$, the overall training objective $\ell_{\text{total}}$ is then a weighted sum of these losses:
$\ell_{\text{total}}=\ell_{\text{cmd}}+\gamma_{1}\,\ell_{\text{cls}}+\gamma_{2}\,\ell_{\text{tgt}},$ (10)
@@ 5 Experiments and results
All data processing and experiments were conducted on a workstation with a 48GB Quadro RTX 8000 GPU and 250GB RAM. We randomly sample 85% of the sessions from the final dataset as the training set and 15% as the validation set (resulting in 1,031,838 test samples), while ensuring that both sets contain all unique commands. The session (sequence) lengths in both sets vary dynamically from 5 to 110 to best simulate real-world production environments. The validation follows the standard approach in sequential recommendation: only the last command in the input sequence is masked, and the model’s performance is evaluated based on its ability to predict the masked item given the preceding context.
@@ 5.2 Comparison of Transformer architectures
Overall, our proposed model architecture outperforms the baselines across all metrics, regardless of the Transformer backbone used. This also demonstrates that the augmented features introduced in Section 3.1.4 have a generalizable positive effect on improving command recommendation. Comparing different backbone networks, decoder-only architectures such as Mixtral-MoE and Llama3.2 outperform BERT and T5 due to their more recent designs and advanced techniques, despite the latter two having the ability to learn bidirectional contextual information. Among them, Mixtral-MoE achieves the best performance, with the correct next-step command appearing in the top-5 and top-10 recommendations with approximately 71.7% and 83.3% probability, respectively.
@@ 7.1 Dataset scale and comparison with previous studies
The dataset used in this study surpasses the scale of previous research related to design command prediction. Pan et al. [3] used BIM logs from an architecture company spanning roughly two years, with a total of 352,056 lines and 14 classes of commands to be predicted. Gao et al. [2] collected BIM logs from a university course on learning 3D modeling software, yielding 490,462 lines with 89 different commands. Matejka et al. from Autodesk [21] collected around 40,000,000 interaction tuples from 16,000 AutoCAD users over six months through the Customer Involvement Program, and used a statistical collaborative filtering algorithm to recommend the next command. In comparison, our final dataset comprises 375,304,719 lines and 4,939 kinds of commands collected in six months. These logs capture anonymous user operations in Vectorworks across different countries, disciplines, and projects, going beyond purely modeling commands to include rendering, lighting, simulation, and more. Such breadth demands more comprehensive and efficient data processing methods and more complex neural network designs. This is also one of our key motivations for adapting modern deep sequential recommendation systems to BIM command prediction.
@@ 9 Data availability statement
The code will be open-sourced soon. The dataset used in this study is confidential.
@@ Appendix A Detailed information of the dataset used in this study
Rectangle; Extrude and Edit Workflow Object The command entry indicates the utilization of the Rectangle tool in Vectorworks, which supports various creation modes including Corner to Corner and Push/Pull for 3D extrusion. The ”End Event: Change Attributes” signifies the completion of an operation that alters the visual properties of selected objects, impacting their appearance and ensuring design consistency within the project.

### v2

@@ Abstract
The adoption of Building Information Modeling (BIM) and model-based design within the Architecture, Engineering, and Construction (AEC) industry has been hindered by the perception that using BIM authoring tools demands more effort than conventional 2D drafting. To enhance design efficiency, this paper proposes a BIM command recommendation framework that predicts the optimal next actions in real-time based on users’ historical interactions. We propose a comprehensive filtering and enhancement method for large-scale raw BIM log data and introduce a novel command recommendation model. Our model builds upon the state-of-the-art Transformer backbones originally developed for large language models (LLMs), incorporating a custom feature fusion module, dedicated loss function, and targeted learning strategy. In a case study, the proposed method is applied to over 32 billion rows of real-world log data collected globally from the BIM authoring software Vectorworks. Experimental results demonstrate that our method can learn universal and generalizable modeling patterns from anonymous user interaction sequences across different countries, disciplines, and projects. When generating recommendations for the next command, our approach achieves a Recall@10 of approximately 84%. The code is available at: https://github.com/dcy0577/BIM-Command-Recommendation.git
@@
[…]
@@ 2.2 Transformers for sequential recommendation
Recent advancements in large language models (LLMs) have demonstrated remarkable capabilities in modeling sequential data through the Transformer architecture. Originally developed for Natural Language Processing (NLP), Transformers have been successfully adapted to recommendation systems due to their ability to capture complex dependencies in sequential user interactions [22, 23, 24]. The attention mechanism enables effective modeling of user behavior dynamics, making it well-suited for scenarios where user actions and interests evolve across sessions [1].
While NLP Transformers typically process text through tokenizers and task-specific prediction heads, recommendation systems adapt this architecture by incorporating meta-information such as user comments, product descriptions, and prices in e-commerce domains [25]. Our study leverages the core Transformer architecture to model sequential dependencies among commands, replacing NLP-specific components with a custom feature fusion module and a prediction head tailored for command recommendation.
Our implementation builds on Transformer4Rec [24], an open-source framework that bridges NLP and sequential recommendation systems using Transformer backbones from models like XLNet [26]. This framework has demonstrated state-of-the-art performance in news and e-commerce domains. However, our research explores BIM command recommendation, which presents unique challenges including limited meta-information, varying session lengths, and long-tail command distributions.
@@ 3.1 Data filtering and enhancement
Previous studies have primarily focused on developing data processing pipelines based on Revit logs; a comprehensive review is provided in [5]. However, these studies have not addressed more challenging issues related to logs from other BIM authoring software, which often exhibit multilingual content, a mix of high- and low-level commands, and excessive complexity. As an example, Figure 3 illustrates a snippet of the Vectorworks log containing two sessions in different languages.
@@ 3.1.2 Multi-language alignment
If the similarity falls below the threshold, we assume that the ID is shared by multiple different commands, as the English names obtained by the translation API show significant semantic differences. In such cases, the unsupervised clustering algorithm DBSCAN [28, 29] is employed to further group the commands into sub-clusters. Each sub-cluster is subsequently labeled using the command name of its midpoint vector. The process is outlined in Alg. 2. The example outcomes from this module are illustrated in Table 1.
@@ 3.2.5 Multi-task learning and loss function
where $p_{t}$ is the model’s estimated probability for the ground-truth class, $\alpha_{t}$ is a weighting factor to address class imbalance, and $\gamma$ is a focusing parameter that smoothly adjusts the rate at which easy examples are down-weighted. Let the losses for the three tasks be $\ell_{\text{cmd}},\ell_{\text{typ}},\ell_{\text{tgt}}$, the overall training objective $\ell_{\text{total}}$ is then a weighted sum of these losses:
$\ell_{\text{total}}=\ell_{\text{cmd}}+\gamma_{1}\,\ell_{\text{typ}}+\gamma_{2}\,\ell_{\text{tgt}},$ (10)
@@ 5 Experiments and results
All data processing and experiments were conducted on a workstation with a 48GB Quadro RTX 8000 GPU and 250GB RAM. We randomly sample 85% of the sessions from the final dataset as the training set and 15% as the validation set (resulting in 1,031,839 test samples), while ensuring that both sets contain all unique commands. The session (sequence) lengths in both sets vary dynamically from 5 to 110 to best simulate real-world production environments. The validation follows the standard approach in sequential recommendation: only the last command in the input sequence is masked, and the model’s performance is evaluated based on its ability to predict the masked item given the preceding context.
@@ 5.2 Comparison of Transformer architectures
Overall, our proposed model architecture outperforms the baselines across all metrics, regardless of the Transformer backbone used. We further employed the paired bootstrap method to compare the performance of our proposed model with the baseline. The calculated 95% confidence intervals and p-values demonstrate that the improvement is statistically significant. Detailed results can be found in D. This also demonstrates that the augmented features introduced in Section 3.1.4 have a generalizable positive effect on improving command recommendation. Comparing different backbone networks, decoder-only architectures such as Mixtral-MoE and Llama3.2 outperform BERT and T5 due to their more recent designs and advanced techniques, despite the latter two having the ability to learn bidirectional contextual information. Among them, Mixtral-MoE achieves the best performance, with the correct next-step command appearing in the top-5 and top-10 recommendations with approximately 71.7% and 83.3% probability, respectively.
@@ 7.1 Dataset scale and comparison with previous studies
The dataset used in this study surpasses the scale of previous research related to design command prediction. Pan et al. [3] used BIM logs from an architecture company spanning roughly two years, with a total of 352,056 lines and 14 classes of commands to be predicted. Gao et al. [2] collected BIM logs from a university course on learning 3D modeling software, yielding 490,462 lines with 89 different commands. Matejka et al. from Autodesk [21] collected around 40,000,000 interaction tuples from 16,000 AutoCAD users over six months through the Customer Involvement Program, and used a statistical collaborative filtering algorithm to recommend the next command. In comparison, our final dataset comprises 753,800,816 lines and 4,939 kinds of commands collected in six months. These logs capture anonymous user operations in Vectorworks across different countries, disciplines, and projects, going beyond purely modeling commands to include rendering, lighting, simulation, and more. Such breadth demands more comprehensive and efficient data processing methods and more complex neural network designs. This is also one of our key motivations for adapting modern deep sequential recommendation systems to BIM command prediction.
@@
There might be the concern that inappropriate work sequences may get captured and negatively influence the system. Although individual users, even experienced ones, may sometimes perform suboptimal workflows, it is highly improbable that inappropriate or flawed workflows would be systematically repeated by such a large and diverse user base. Instead, effective and common command sequences will emerge as strong signals, while bad or erroneous sequences will fall into statistical noise and are unlikely to be learned by the model as general patterns.
@@ 8 Conclusion
The conceptualization, aesthetics, and functionality of architecture involve highly personalized and complex design decisions. The proposed system does not attempt to automate such high-level design decisions, or replace the human designers. We emphasize that the quality of the final building design still depends on the designer’s knowledge and decisions, while the proposed system supports the software modeling process needed to realize those design intents. Our approach aims to improve user efficiency, reduce modeling time, and lower the learning curve.
@@ 9 Data availability statement
The code is available at: https://github.com/dcy0577/BIM-Command-Recommendation.git. The dataset used in this study is confidential.
@@ Appendix A Detailed information of the dataset used in this study
Rectangle; Extrude and Edit Workflow Object The log indicates that the user has utilized the Rectangle tool within Vectorworks, specifically accessing the Extrude and Edit menu to convert a 2D rectangle into a 3D form. This operation allows for the specification of height, enabling the user to create complex geometries from simple shapes while leveraging the Push/Pull mode for immediate extrusion in 3D views. Overall, this workflow enhances the modeling capabilities of the software by facilitating the transformation of basic 2D elements into detailed 3D objects.
@@
Appendix B Data processing algorithms
Algorithm 1 Matching and Removal of Redo and Undo Commands
1:Initialize $recent\_commands\leftarrow\emptyset$
2:Initialize $recent\_undone\_commands\leftarrow\emptyset$
3:Initialize $to\_remove\_indices\leftarrow\emptyset$
4:for each $(\text{cmd},idx)$ in each log session do
5:    if cmd is a normal command then
6:        Append $(\text{cmd},idx)$ to $recent\_commands$
[…]
8:        Search backward in $recent\_commands$ for a matching command $(c,i)$
9:        if found then
10:           Add $idx$ and $i$ to $to\_remove\_indices$
11:           Remove $(c,i)$ from $recent\_commands$
12:           Append $(c,i)$ to $recent\_undone\_commands$
13:        end if
[…]
16:        if found then
17:           Remove $i$ from $to\_remove\_indices$
18:           Add $idx$ to $to\_remove\_indices$
19:           Remove $(c,i)$ from $recent\_undone\_commands$
20:           Append $(c,i)$ to $recent\_commands$
21:        else
22:           Add $idx$ to $to\_remove\_indices$
23:        end if
24:    end if
25:end for
Algorithm 2 Multi-language Alignment
1:Actual modeling logs S
2:Predefined similarity threshold $\phi$
3:Aligned logs $\textbf{S}_{\textit{aligned}}$
6:for $\textbf{S}_{\textit{id}}\in\textbf{S}_{\textit{id\_groups}}$ do
7:    for $\text{s}\in\textbf{S}_{\textit{id}}$ do
10:    end for
12:    if $\phi_{\textit{id}}\geq\phi$ then
15:    else
20:        end for
21:    end if
22:end for
25:return $\textbf{S}_{\textit{aligned}}$
Algorithm 3 Redundant Command Identification
1:Aligned logs $\textbf{S}_{\textit{aligned}}$
2:Similarity threshold $\phi$
3:High-/low-level command mapping $\textbf{S}_{\textit{mapping}}$
4:$\textbf{S}_{\textit{mapping}}\leftarrow\emptyset$
7:    $\textbf{S}_{\textit{pairs}}\leftarrow\emptyset$
[…]
13:        if $\text{confidence}>\phi$ then
15:           if approved then
17:           end if
18:        end if
19:    end for
20:    if $\textbf{S}_{\textit{pairs}}\neq\emptyset$ then
22:    end if
23:end for
24:return $\textbf{S}_{\textit{mapping}}$
Appendix C Hyperparameter settings in data filtering and enhancement
In this section, we explain the rationale behind our settings of two hyperparameter values in the data filtering and enhancement pipeline: the similarity threshold for multi-language alignment, as described in Section 3.1.2, and the confidence threshold for association rule mining (ARM), as discussed in Section 3.1.3.
In multi-language alignment, if the similarity threshold is too low, we may miss cases where the same ID corresponds to semantically different commands. On the other hand, if the threshold is too high, unsupervised clustering may incorrectly separate semantically similar commands into different clusters due to slight variations in translation. For instance, in ID 424, the translated versions contain five instances of End Event: Edit Subdivision (424) and one instance of End Event: Subdivision Change (424). However, the command End Event: Subdivision Change (424) may be clustered into a different group, despite referring to the same semantic action.
We tested five different similarity thresholds: 0.95, 0.90, 0.85, 0.80, and 0.75. For each threshold, command IDs with similarity above the threshold were directly labeled, while those below the threshold were processed using unsupervised clustering. To evaluate the quality of this process, we performed stratified sampling [60] by selecting 10% of the IDs from both the above-threshold and below-threshold groups, due to the large size of the dataset. We then manually inspected the sampled results and calculated the error rate. Errors were defined as either (1) incorrect clustering of semantically different commands or (2) misclustering of semantically similar commands. The result can be shown in Figure 19.
Figure 19: Sensitivity analysis of similarity threshold
The error rate is lowest around the threshold of 0.80. In this study, we applied a slight buffer and set the threshold to 0.82. However, the overall error rate remained relatively high, at approximately 30%. To further enhance accuracy, we conducted an additional manual verification step. This involved reviewing and refining the command dictionary by checking whether the translations were accurate and semantically aligned.
The confidence in ARM measures the strength of the relationship between the two commands in a command pair. A higher confidence value indicates a stronger association: when the first command occurs, the corresponding command in the pair is highly likely to be triggered afterward. For example, the confidence of the command pair Tool: Line (-201) and End Event: Create Line (166) is 0.93, indicating a strong association between them.
However, due to the large size of the dataset, a single command can form a large number of command pairs. Additionally, some commands do not trigger any follow-up commands. For example, Menu: Close (-4) forms around 40,000 command pairs after ARM, yet none of them have a confidence value greater than 0.4. We further tested this command in the software environment and confirmed that it does not trigger any follow-up commands. Thus, to reduce the workload caused by the abundance of obviously irrelevant command pairs, we applied a confidence threshold to automatically filter them out. Specifically, we set a relatively low threshold of 0.4, which effectively eliminates the most apparent non-triggering relationships and significantly reduces the need for manual verification. After this filtering step, we further conducted manual refinement within the software to validate and improve the remaining high-confidence command pairs.
Overall, the settings of these two hyperparameters have minimal impact on the final processed data, as all results are subject to additional manual review.
Appendix D Statistical significance test results between Proposed and Baseline models
Table 8: Statistical significance test results between Proposed and Baseline models using paired bootstrap [%]
Backbone Recall@k NDCG@k k=3 k=5 k=10 k=3 k=5 k=10 BERT (Encoder-only) Mean Difference 0.3714 0.3043 0.1026 0.4119 0.3881 0.3225 95% CI [0.3191, 0.4244] [0.2577, 0.3520] [0.0633, 0.1413] [0.3728, 0.4506] [0.3551, 0.4206] [0.2944, 0.3509] p-value (two-sided) $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ Mixtral-MoE (Decoder-only) Mean Difference 1.2249 0.9716 0.5847 1.2646 1.1610 1.0342 95% CI [1.1659, 1.2825] [0.9204, 1.0220] [0.5436, 0.6253] [1.2210, 1.3078] [1.1239, 1.1966] [1.0029, 1.0660] p-value (two-sided) $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ Llama3.2 (Decoder-only) Mean Difference 1.2533 0.8964 0.6013 1.2226 1.0758 0.9800 95% CI [1.1944, 1.3118] [0.8445, 0.9489] [0.5597, 0.6425] [1.1782, 1.2675] [1.0377, 1.1137] [0.9473, 1.0126] p-value (two-sided) $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ T5 (Encoder-Decoder) Mean Difference 0.4802 0.3756 0.2446 0.4920 0.4497 0.4064 95% CI [0.4271, 0.5339] [0.3292, 0.4220] [0.2066, 0.2825] [0.4540, 0.5313] [0.4164, 0.4821] [0.3783, 0.4348] p-value (two-sided) $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
- Mean Difference: Average performance improvement of Proposed over Baseline model (Proposed - Baseline) across 10,000 bootstrap samples
- 95% CI: 95% confidence interval from 10,000 bootstrap iterations
- p-value computed via permutation test [61]
To assess whether the performance differences between our proposed models and the baselines are statistically significant, we employed paired bootstrap to calculate the confidence intervals and paired permutation tests to obtain p-values.
Bootstrap confidence intervals. For each evaluation metric, we: (1) Calculated the paired differences between the two models: $\Delta_{i}=X_{1i}-X_{2i}$, where $X_{1i}$ and $X_{2i}$ represent the metric values for the $i$-th test sample from the proposed and baseline models, respectively. (2) Generated 10,000 bootstrap samples by resampling the differences with replacement. (3) Computed the mean difference for each bootstrap sample and derived the 95% confidence interval using the 2.5th and 97.5th percentiles of the bootstrap distribution.
Paired permutation test. To obtain p-values, we conducted paired permutation tests [61] that account for the paired nature of our comparisons (same test samples evaluated by both models). The procedure is as follows: (1) Computed the observed test statistic: $T_{\text{obs}}=\text{mean}(X_{1})-\text{mean}(X_{2})$ (2) Under the null hypothesis of no difference between models, we randomly exchanged the model assignments within each pair with probability 0.5, repeated 10,000 times. (3) For each permutation $\pi$, calculated the test statistic: $T_{\pi}=\text{mean}(X_{1\pi})-\text{mean}(X_{2\pi})$ (4) Computed the two-sided p-value as: $p=P(|T_{\pi}|\geq|T_{\text{obs}}|)$
The detailed results in Table 8 demonstrate that the improvements of the proposed models over the baselines are statistically significant.